A boat crosses a river that flows downstream at a rate of 11 m/s. The boat travels at a
speed of 17 m/s.
a. If the boat wishes to dock directly across the river from where it departs, at what
angle must it steer to make it to the dock?
b. What is the resultant speed of the boat?
1
Expert's answer
2020-12-02T09:50:49-0500

a)


"\alpha=\arcsin{\frac{11}{17}}=40\degree"

b)


"v=\sqrt{17^2-11^2}=13\frac{m}{s}"